# CrowdVLM-R1: Expanding R1 Ability to Vision Language Model for Crowd Counting using Fuzzy Group Relative Policy Reward

### v1

@@ Abstract
We propose Fuzzy Group Relative Policy Reward (FGRPR), a novel framework that integrates Group Relative Policy Optimization (GRPO) with a fuzzy reward function to enhance learning efficiency. Unlike the conventional binary 0/1 accuracy reward, our fuzzy reward model provides nuanced incentives, encouraging more precise outputs. Experimental results demonstrate that GRPO with a standard 0/1 accuracy reward underperforms compared to supervised fine-tuning (SFT). In contrast, FGRPR, applied to Qwen2.5-VL-(3B/7B), surpasses all baseline models—including GPT-4o, LLaMA2-90B, and SFT—across five in-domain datasets. On an out-of-domain dataset, FGRPR achieves performance comparable to SFT but excels when target values are larger, as its fuzzy reward function assigns higher rewards to closer approximations. This approach is broadly applicable to tasks where the precision of the answer is critical. Code and data: https://github.com/yeyimilk/CrowdVLM-R1
Index Terms:
Vision language models, large language model, crowd counting, DeepSeek R1, reinforcement learning, group relative policy optimization, fuzzy group relative policy reward, crowd counting.
Figure 1: An overview of our framework which integrates group relative policy optimization (GRPO) and the proposed fuzzy group relative policy reward (FGRPR) to train a visual language model (VLM) for crowd counting tasks. From left to right, an image and prompt are input to a base LLM model to obtain initial counting outcomes (The ground truth value on the top left corner of the input image is just an indicator number for visualization purpose. It is only used for rewards calculation but will not send to LLM as part of input). The instruction model uses a policy model, parameterized by $\theta$, to obtain policy and generate outputs. The proposed fuzzy group relative policy reward is then used to adjust the reward for individual image. The adjusted reward is used to calculate the objective function value, which in term update the $\theta$ parameters for supervised fine tuning.
@@ I Introduction
While DeepSeek R1 is an LLM, its training paradigm can be extended to vision-language models (VLMs). There are various types of visual tasks, such as crowd counting [3, 4, 5], image classification [6, 7, 8], and visual question answering [9, 10]; in the context of VLM, they require the model to go beyond simple pattern recognition and perform precise reasoning. Latest works explore applying RL for visual tasks, one of which is VLM-R1 [11], which assigns rewards based on the intersection over union (IoU) score - a binary reward of 1 if IoU $>$ 0.5, and 0 otherwise. However, this approach does not distinguish between near-perfect and marginally correct detections (e.g., an IoU of 0.55 vs. 0.95), then better results do not receive higher rewards. This suggests the need for more fine-grained reward schemes when applying RL to vision-based reasoning tasks.
@@
To address RQ1, we first conduct a systematic investigation into the impact of RL on VLMs. Specifically, we adapt GRPO, originally used in DeepSeek R1, for fine-tuning mainstream VLMs. Our experiments confirm that GRPO is highly feasible for the counting task in VLMs. However, it requires customized reward function, as conventional reward function has certain limitations.
Issues like this highlights the need to address our RQ2. A well-structured reward function should provide feedback with finer granularity and ensure that slight improvements in counting accuracy are properly incentivized rather than treated as binary outcomes. We solve this research question by introducing a fuzzy group relative policy reward (FGRPR). Our reward function considers both the model’s output format and the numerical precision of its predictions. The format reward is designed as a binary signal to encourage the model to provide an actual count for the scene. Meanwhile, the accuracy reward is based on the difference between the predicted count and the ground truth, guiding the model toward more precise predictions by penalizing deviations from the correct count.
@@ II Related Work
Traditional methods for Crowd Counting. Early research on Crowd Counting can be grouped into three categories: detection-based, regression-based and density estimation-based methods. Detection-based approaches typically utilize detectors for a specific object, e.g., human head or animal body, and perform counting based on the detection results[12, 13, 14]. These methods have achieved good accuracy for sparse scenes, but struggle under dense or occluded scenarios. Regression-based methods, on the other hand, maps the image patches directly to the count of objects via regression techniques[15, 16, 17]. Density estimation methods get rid of learning to localize individual objects by estimating the count with density maps [18, 19, 20].
@@
With the success of Convolution Neural Networks (CNNs) in multiple computer vision tasks, researchers start to leverage CNNs to learning non-linear mappings from crowd images to counts. CNNs have been proven to extract features of higher quality than traditional methods for many vision tasks including counting [21]. Among numerous CNN-based approaches, early works such as [22, 23] use CNN models to perform end-to-end regression without significant modification of the model architecture. Subsequent research has focused on enhancing model architectures[24, 25, 26, 27, 28, 29]. Examples include multi-stage CNNs [28], which are designed to adapt to counting scenes with different scales and distortion, and Mixture of CNNs [29] with a mixture-of-exptert approach where each CNNs is specialized to a different scene.
@@
Vision-Language Models (VLMs) for Counting. The need for supervision and labeling of datasets limits the generalizability of both CNN-based and non-CNN-based methods. Recent vision-language pretraining[30][31] stop relying on supervised training by learning from an extremely large number of image-text pairs. VLMs generates high quality visual-textual features, which are adaptable to a broad spectrum of open-world downstream tasks. Recent VLM-based counting methods have demonstrated superior performance compared to class-specific approaches in counting accuracy[32, 33, 34, 35, 36, 37]. These approaches establish a joint embedding space that connects images with textual inputs, hence enable models to comprehend general concepts learned during extensive pretraining.
@@
Reasoning enhancement for Large Language Models. As LLM’s capabilities advance, there is a growing research focus on their reasoning abilities. A recent work, DeepSeek R1 [1], utilizes pure reinforcement learning with rule-based reward and demonstrates strong performance on a series of reasoning tasks. This approach can be naturally extended to the multimodality setting and applied to VLMs [38, 39, 40, 41].
Figure 2: Examples of model outputs at during the training stage, showing various counting strategies. In the left image, where the crowd is dense and difficult to count, the model segments the scene into six regions and estimates the number of individuals in each. In the middle image, with only a few people, the model counts them individually. In the right image, which has a moderate density of sheep, the model scans row by row to ensure each entity is counted once. These strategies resemble human approaches to counting in different scenarios.
@@
In this section we present our approach to improving VLM-based counting through reinforcement learning. We begin by introducing Group Relative Policy Optimization (GRPO) and highlighting its advantages over previous RL methods. Next, we propose a novel fuzzy group relative policy reward (FGRPR), which refines the reward structure by accounting for both output format adherence and numerical precision. Finally, we outline our overall framework, which integrates GRPO and FGRPR into the training pipeline of our proposed model, CrowdVLM-R1, to improve its counting performance across diverse datasets and real-world scenarios.
@@ III-A Group Relative Policy Optimization
The GRPO[2] is an RL algorithm derived as a variant of Proximal Policy Optimization (PPO) [42]. This algorithm was introduced by the Deepseek group, extending the PPO framework to incorporate group-based optimization strategies. The value function used in PPO is typically implemented as a separate model of comparable size to the policy model, introducing significant memory and computational overhead. By utilizing the average reward of multiple sampled outputs generated in response to the same input as the baseline, GRPO eliminates the need for an additional value function approximation. This approach reduces both computational complexity and memory overhead, while maintaining effective variance reduction in policy gradient updates.
For standardized GRPO, given an input question $q$, a group of $G$ outputs $\mathbf{o}=\{o_{1},o_{2},\dots,o_{G}\}$ are generated from a sampling process determined by the old policy $\pi_{\theta_{\text{old}}}$, parameterized by $\theta$, to answer the question $q$. Let $\textbf{r}=\{r_{1},r_{2},\dots,r_{G}\}$ denotes $G$ rewards in the group and $\bar{\textbf{r}}$ is the group average reward, the advantages $\hat{A}_{i,t}$ of all tokens in the output as the normalized reward will be
@@
where $\epsilon$ and $\beta$ are hyperparameters, and $\hat{A}_{i,t}$ represent the advantage computed based on the relative rewards of outputs within each group.$\pi_{\theta}$ is the current policy to be optimized, generated by the trainable policy model, typically an LLM or sequence model, parameterized by $\theta$. $\pi_{old}$ is the behavior policy used for sampling, which comes from a frozen model of the policy $\pi_{\theta}$ from a previous iteration. $\pi_{ref}$ is the reference policy for KL penalty, which is typically from the SFT model (also a LLM or sequence model), trained with human data.
@@
PPO employs a clipped objective function to limit the magnitude of policy updates. However, this clipping mechanism can sometimes overly restrict updates, leading to slower convergence. On the other hand, TRPO ensures the stability of policy updates through a strict KL divergence constraint, but its optimization process requires second-order methods (e.g., conjugate gradient), resulting in high computational costs that make it unsuitable for large-scale reinforcement learning tasks. GRPO addresses these limitations by combining clipping mechanisms with a KL divergence term, ensuring a balance between exploration and stability. This balance makes GRPO more reliable in complex tasks, as it maintains both stability and consistency during optimization. While PPO’s fixed clipping range may limit the convergence speed of policies, GRPO introduces a dynamic adjustment mechanism (e.g., through a beta parameter) to adaptively control the policy update range, effectively adjusting the KL divergence to achieve stable and efficient policy updates.
In previous work like VLM-R1 [11], a typical GPRO accuracy reward is defined as
@@
where $\hat{y}$ is the count predicted by the model and $y$ denotes the real count. This is a ”fixed reward” that assigns 1 to any prediction values lie in the range $(\frac{y}{2},\frac{3y}{2})$. For example, if the image contains 500 persons, the prediction values of 260, 500, 740 share the same reward and have no difference from model perspective. However, these values are supposed to have different rewards for their different accuracies.
@@ III-B FGRPR: Fuzzy Group Relative Policy Reward
To enable SFT by using multiple sources of crowd counting data, we introduce FGRPR method, which employs a fuzzy reward function to gauge the fine-tuning of the R1 model by taking qualify of counting, compared to the benchmark, for consideration.
Overall, the reward function consists of two parts: format reward, $r_{f}$, and precision reward, $r_{p}$ which are linearly combined in Eq. (4).
@@
The format reward is straightforward, that is, if the model output format($\hat{y}_{p}$) , follows the defined format($\mathbb{F}$), $r_{f}$ is $1$, otherwise, $r_{f}$ is 00.
@@
Previous work and research topics indicated that the answers can be or simply categorized as “Yes” or “No” which means the precision reward can be easily used with binary value, “1” or “0”. While the crowd counting task also has a “ground true” value, we prefer to have a fuzzy reward to represent the model estimated value that it should have a higher reward when the answer is closer to number of objects and lower reward or even 0 if the predicted value is far away from the “ground truth” value. Thus, the precision reward in our task is defined as shown below.
@@ III-C Framework
Figure 3: A conceptual view of R1 model expansion to vision language model for crowd counting (CrowdVLM+R1) using fuzzy group relative policy reward. Multiple crowd-sourced inputs, with ground-truth counting numbers, are used for supervised fine tuning. The inputs are fed into to a base model to generate outcomes, followed by the proposed fuzzy group relative policy reward (FGRPR) to determine individual rewards and system objective function value $\mathcal{J}_{GRPO}(\theta)$. Gradient propagtation process is then applied to update $\theta$, and in terms improve the base model for better counting results.
Figure 4: This figure presents the image examples from test dataset where each column comes from the same original dataset and the first row shows an easier counting task while the second one is more complex within their original data sources.
@@
Dataset Type Sheep Video Game Characters Street Pedestrians Wheat Heads Cars Manatees*
@@
    Out-of-domain dataset.
Fig. 3 presents the general idea of training a CrowdVLM-R1 model. Instead of using a labeled dataset to fine-tune the based instruct model to optimize its performance directly, in our method, GRPO is employed during the SFT stage with verifiable rewards (Equ. 4, including format reward (Equ. 6) and counting precision reward (Equ. 7).
For details, our proposed framework is shown in Fig. 1. A structured prompt comprises a natural language description (including the counting task and the expected output format) and an input image. This prompt is input into a policy model—implemented as a LLM—which defines the learner policy $\pi_{\theta}$ and autoregressively generates a set of candidate outputs $\{o_{1},o_{2},\ldots,o_{G}\}$. Each output $o_{i}$ consists of a step-by-step reasoning process and a predicted count corresponding to the target objects in the input image. Subsequently, a reward model evaluates these generated responses, which assigns a scalar reward $r_{i}$ to each output. The reward $r_{i}$ is composed of two components: a format reward $r_{f_{i}}$ and a prediction reward $r_{p_{i}}$, as detailed in Section III-B. These individual rewards are then aggregated to compute group-wise advantages $A_{i}$ for all $G$ outputs, reflecting the relative quality of each sampled response. Our objective function $\mathcal{J}_{\mathrm{GRPO}}(\theta)$ is formulated using the computed advantages $A_{i}$ (for $i=1,\ldots,G$) and a KL-divergence regularization term $D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$, scaled by a hyperparameter $\beta$ in Equation (2). Here, in our case, the reference model is an SFT LLM model trained with human data. We will minimize the objective/loss function using gradient ascent (or descent on the negative objective). After the calculation, the gradient signal will flow from the objective back into the policy model to update its parameters $\theta$ based on both the relative advantages and the regularization signal. After sufficient interactions, we will finally receive a trained-well LLM model that can count the number of specified targets in a given image.
@@
Model In-Domain Out-Domain Overall
@@
  • Q: Qwen-VL-2.5-Instruct; R1: model was trained with GRPO; FR: model was trained with FGRPR;

### v2

@@ Abstract
We propose CrowdVLM-R1, which expands the R1 base model for accurate crowd counting, using a novel framework that integrates the fuzzy group relative policy optimization reward function (FGRPR) to enhance learning efficiency. Unlike the conventional binary (0/1) accuracy reward, our fuzzy reward model, FGRPR, which contains both format and precision rewards, provides nuanced incentives to encourage the R1 model to learn to adjust policies towards precise outputs. Supervised fine-tuning (SFT) is also integrated for the CrowdVLM-R1 model to learn from a handful of inputs to enable both in-domain and out-of-domain counting. Experimental results demonstrate that GRPO with a standard binary accuracy reward underperforms compared to SFT. In contrast, FGRPR, applied to Qwen2.5-VL-(3B/7B), surpasses all baseline models, including GPT-4o, LLaMA2-70B and SFT, in five domain datasets. For out-of-domain datasets, FGRPR achieves performance comparable to SFT but excels when target values are larger, as its fuzzy reward function assigns higher rewards to closer approximations. This approach is broadly applicable to tasks where the precision of the answer is critical. The code and data are available at: https://github.com/yeyimilk/CrowdVLM-R1
Figure 1: An overview of our framework, which integrates group relative policy optimization (GRPO) and the proposed fuzzy group relative policy reward (FGRPR) to train a visual language model (VLM) for crowd counting tasks. From left to right, an image and prompt are input to a base LLM to obtain initial counting outcomes (The ground truth value on the top left corner of the input image is just an indicator number for visualization purposes. It is only used for rewards calculation, but will not send to LLM as part of the input. The instruction model uses a policy model, parameterized by $\theta$, to obtain policy and generate outputs. The proposed fuzzy group relative policy reward is then used to adjust the reward for the individual image. The adjusted reward is used to calculate the objective function value, which in turn updates the $\theta$ parameters for supervised fine-tuning.
@@ I Introduction
While DeepSeek R1 is an LLM, its training paradigm can be extended to vision-language models (VLMs). There are various types of visual tasks, such as crowd counting [3, 4, 5], image classification [6, 7, 8], and visual question answering [9, 10]; in the context of VLM, they require the model to go beyond simple pattern recognition and perform precise reasoning. Latest works explore applying RL for visual tasks, one of which is VLM-R1 [11], which assigns rewards based on the intersection over union (IoU) score - a binary reward of 1 if IoU >> 0.5, and 0 otherwise. However, this approach does not distinguish between near-correct and marginally correct detections (e.g., an IoU of 0.55 vs. 0.95), then better results do not receive higher rewards. This suggests the need for more fine-grained reward schemes when applying RL to vision-based reasoning tasks.
@@
To address RQ1, we first conduct a systematic investigation into the impact of RL on VLMs. Specifically, we adapt GRPO, originally used in DeepSeek R1, for fine-tuning mainstream VLMs. Our experiments confirm that GRPO is highly feasible for the counting task in VLMs. However, it requires a customized reward function, as conventional reward function has certain limitations.
Issues like this highlight the need to address our RQ2. A well-structured reward function should provide feedback with finer granularity and ensure that slight improvements in counting accuracy are properly incentivized rather than treated as binary outcomes. We solve this research question by introducing a fuzzy group relative policy reward (FGRPR). Our reward function considers both the model’s output format and the numerical precision of its predictions. The format reward is designed as a binary signal to encourage the model to provide an actual count for the scene. Meanwhile, the accuracy reward is based on the difference between the predicted count and the ground truth, guiding the model toward more precise predictions by penalizing deviations from the correct count.
@@ II Related Work
Traditional methods for Crowd Counting. Early research on Crowd Counting can be grouped into three categories: detection-based, regression-based, and density estimation-based methods. Detection-based approaches typically utilize detectors for a specific object, e.g., human head or animal body, and perform counting based on the detection results [12, 13]. These methods have achieved good accuracy for sparse scenes, but struggle under dense or occluded scenarios. Regression-based methods, on the other hand, map the image patches directly to the count of objects via regression techniques [14]. Density estimation methods get rid of learning to localize individual objects by estimating the count with density maps [15, 16].
@@
With the success of Convolution Neural Networks (CNNs) in multiple computer vision tasks, researchers have started to leverage CNNs to learning non-linear mappings from crowd images to counts. CNNs have been proven to extract features of higher quality than traditional methods for many vision tasks, including counting [17]. Among numerous CNN-based approaches, early works such as [18, 19] use CNN models to perform end-to-end regression without significant modification of the model architecture. Subsequent research has focused on enhancing model architectures [20, 21, 22, 23]. Examples include multi-stage CNNs [22], which are designed to adapt to counting scenes with different scales and distortion, and Mixture of CNNs [23] with a mixture-of-exptert approach where each CNNs is specialized to a different scene.
@@
Vision-Language Models (VLMs) for Counting. The need for supervision and labeling of datasets limits the generalizability of both CNN-based and non-CNN-based methods. Recent vision-language pretraining works [24][25] do not rely on supervised training by learning from an extremely large number of image-text pairs. VLMs generates high-quality visual-textual features, which are adaptable to a broad spectrum of open-world downstream tasks. Recent VLM-based counting methods have demonstrated superior performance compared to class-specific approaches in counting accuracy [26, 27, 28, 29, 30, 31]. These approaches establish a joint embedding space that connects images with textual inputs, and hence enable models to comprehend general concepts learned during extensive pretraining.
@@
Reasoning enhancement for Large Language Models. As LLMs’ capabilities advance, there is a growing research focus on their reasoning abilities. A recent work, DeepSeek R1 [1], utilizes pure reinforcement learning with rule-based reward and demonstrates strong performance on a series of reasoning tasks. This approach can be naturally extended to the multimodality setting and applied to VLMs [32, 33, 34, 35, 36].
@@
Figure 2: Examples of model outputs during the training stage, showing various counting strategies. In the left image, where the crowd is dense and difficult to count, the model segments the scene into six regions and estimates the number of individuals in each. In the middle image, with only a few people, the model counts them individually. In the right image, which has a moderate density of sheep, the model scans row by row to ensure each entity is counted once. These strategies resemble human approaches to counting in different scenarios.
In this section, we present our approach to improving VLM-based counting through reinforcement learning. We begin by introducing Group Relative Policy Optimization (GRPO) and highlighting its advantages over previous RL methods. Next, we propose a novel fuzzy group relative policy reward (FGRPR), which refines the reward structure by accounting for both output format adherence and numerical precision. Finally, we outline our overall framework, which integrates GRPO and FGRPR into the training pipeline of our proposed model, CrowdVLM-R1, to improve its counting performance across diverse datasets and real-world scenarios.
@@ III-A Group Relative Policy Optimization
GRPO [2] is an RL algorithm derived as a variant of Proximal Policy Optimization (PPO) [37]. This algorithm was introduced by the Deepseek group, extending the PPO framework to incorporate group-based optimization strategies. The value function used in PPO is typically implemented as a separate model of comparable size to the policy model, introducing significant memory and computational overhead. By utilizing the average reward of multiple sampled outputs generated in response to the same input as the baseline, GRPO eliminates the need for an additional value function approximation. This approach reduces both computational complexity and memory overhead, while maintaining effective variance reduction in policy gradient updates.
For standardized GRPO, given an input question $q$, a group of $G$ outputs $\mathbf{o}=\{o_{1},o_{2},\dots,o_{G}\}$ are generated from a sampling process determined by the old policy $\pi_{\theta_{\text{old}}}$, parameterized by $\theta$, to answer the question $q$. Let $\textbf{r}=\{r_{1},r_{2},\dots,r_{G}\}$ denote $G$ rewards in the group and $\bar{\textbf{r}}$ be the group average reward, the advantages $\hat{A}_{i,t}$ of all tokens in the output as the normalized reward will be
@@
where $\epsilon$ and $\beta$ are hyperparameters, and $\hat{A}_{i,t}$ represents the advantage computed based on the relative rewards of outputs within each group. $\pi_{\theta}$ is the current policy to be optimized, generated by the trainable policy model, typically an LLM or sequence model, parameterized by $\theta$. $\pi_{old}$ is the behavior policy used for sampling, which is obtained from a frozen model of the policy $\pi_{\theta}$ trained in a previous iteration. $\pi_{ref}$ is the reference policy for the KL penalty, which is typically from the SFT model (also a LLM or sequence model).
@@
PPO employs a clipped objective function to limit the magnitude of policy updates. However, this clipping mechanism can sometimes overly restrict updates, leading to slower convergence. On the other hand, Trust Region Policy Optimization (TRPO) ensures the stability of policy updates through a strict KL divergence constraint, but its optimization process requires second-order methods (e.g., conjugate gradient), resulting in high computational costs that make it unsuitable for large-scale reinforcement learning tasks. GRPO addresses these limitations by combining clipping mechanisms with a KL divergence term, ensuring a balance between exploration and stability. This balance makes GRPO more reliable in complex tasks, as it maintains both stability and consistency during optimization. While PPO’s fixed clipping range may limit the convergence speed of policies, GRPO introduces a dynamic adjustment mechanism (e.g., through a beta parameter) to adaptively control the policy update range, effectively adjusting the KL divergence to achieve stable and efficient policy updates.
In previous work like VLM-R1 [11], a typical GRPO accuracy reward is defined as
@@
where $\hat{y}$ is the count predicted by the model and $y$ denotes the real count. This is a ”fixed reward” that assigns 1 to any prediction values lie in the range $(\frac{y}{2},\frac{3y}{2})$. For example, if the image contains $500$ persons, the prediction values of $260$, 500500, and $740$ share the same reward and have no difference from a model perspective. However, these values are supposed to have different rewards for their different accuracies.
@@ III-B FGRPR: Fuzzy Group Relative Policy Reward
To enable SFT by using multiple sources of crowd counting data, we introduce the FGRPR method, which employs a fuzzy reward function to gauge the fine-tuning of the R1 model by taking quality of counting, compared to the benchmark, for consideration. The general reward function is defined as
  • Multi-factor Reward: This reward can describe the task in meaningful dimensions (e.g., format, accuracy, precision, coherence, diversity).
  • Modular Design: The total reward is a weighted sum:
    $r(s,a)=\sum_{i}\mu_{i}(s,a)r_{i}(s,a),$
    where each $r_{i}(s,a)$ evaluates a specific aspect.
For the crowd counting task (specifically in 2 dimensional images), one factor is format adherence, and another is counting precision. Overall, the reward function consists of two parts: format reward, $r_{f}$, and precision reward, $r_{p}$, which are linearly combined in Eq. (4).
@@
The format reward is straightforward: if the model output format ($\hat{y}_{p}$) follows the defined format ($\mathbb{F}$), $r_{f}$ is $1$, otherwise, $r_{f}$ is 0. The membership degree $\mu_{f}(s,a)$ quantifies how closely the output $a$ matches the intended format in state $s$, forming the fuzzy set ”format-compliant” (e.g., if the output closely matches but still deviates, $\mu_{f}$ may be 0.8 rather than a crisp 1). Likewise, $\mu_{p}(s,a)$ represents the fuzzy set ”precision-adequate” (e.g., $\mu_{p}$ = 0.6 if the numeric answer is close but not exact). For all ablation experiments presented, we set both $\mu_{f}$ and $\mu_{p}$ defaults to 1 to equally balance format and precision.
@@
Previous work indicated that the answers can be simply categorized as ”Yes” or ”No”, which means the precision reward can be easily used with binary value—1 or 0. While the crowd counting task also has a ground truth value, we prefer to have a fuzzy reward to represent the model estimated value that it should have a higher reward when the answer is closer to number of objects and lower reward or even 0 if the predicted value is far away from the ground truth value. Thus, the precision reward in our task is defined as shown below.
@@ III-C Framework
Figure 3: A conceptual view of R1 model expansion to vision language model for crowd counting (CrowdVLM+R1) using fuzzy group relative policy reward. Multiple crowd-sourced inputs, with ground-truth counting numbers, are used for supervised fine-tuning. The inputs are fed into a base model to generate outcomes, followed by the proposed fuzzy group relative policy reward (FGRPR) to determine individual rewards and system objective function value $\mathcal{J}_{GRPO}(\theta)$. Gradient propagation process is then applied to update $\theta$, and in terms of improving the base model for better counting results.
Figure 4: This figure presents the image examples from the test dataset where each column comes from the same original dataset and the first row shows an easier counting task while the second one is more complex within their original data sources.
@@
Dataset Type Sheep Video Game Characters Street Pedestrians Wheat Heads Cars Manatees (out-of-domain)
@@
Fig. 3 presents the general idea of training a CrowdVLM-R1 model. Instead of using a labeled dataset to fine-tune the base instruction model to optimize its performance directly, in our method, GRPO is employed during the SFT stage with verifiable rewards (equation 4), including format reward (equation 6) and counting precision reward (equation 7).
For details, our proposed framework is shown in Fig. 1. A structured prompt comprises a natural language description (including the counting task and the expected output format) and an input image. This prompt is the input into a policy model—implemented as a LLM—which defines the learner policy $\pi_{\theta}$ and autoregressively generates a set of candidate outputs $\{o_{1},o_{2},\ldots,o_{G}\}$. Each output $o_{i}$ consists of a step-by-step reasoning process and a predicted count corresponding to the target objects in the input image. Subsequently, a reward model evaluates these generated responses, which assigns a scalar reward $r_{i}$ to each output. The reward $r_{i}$ is composed of two components: a format reward $r_{f_{i}}$ and a prediction reward $r_{p_{i}}$, as detailed in Section III-B. These individual rewards are then aggregated to compute group-wise advantages $A_{i}$ for all $G$ outputs, reflecting the relative quality of each sampled response. Our objective function $\mathcal{J}_{\mathrm{GRPO}}(\theta)$ is formulated using the computed advantages $A_{i}$ (for $i=1,\ldots,G$) and a KL-divergence regularization term $D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$, scaled by a hyperparameter $\beta$ in Equation (2). Here, in our case, the reference model is an SFT LLM trained with human data. We will minimize the objective/loss function using gradient ascent (or descent on the negative objective). After the calculation, the gradient signal will flow from the objective back into the policy model to update its parameters $\theta$ based on both the relative advantages and the regularization signal. After sufficient interactions, we can finally obtain a model that can accurately count the number of specified targets in a given image.
@@
Model In-Domain Out-of-Domain Overall